# Invariant Federated Learning: A Novel Approach to Addressing Challenges in Federated Learning for Edge Intelligence

### v1

@@ Abstract
Federated learning (FL) has become a crucial solution for distributed learning in edge intelligence, addressing communication constraints and privacy protection. However, challenges such as heterogeneous and asynchronous clients significantly impact model performance. This paper analyzes the harm of abnormal clients through parameter orthogonal decomposition innovatively and shows that the exit of abnormal clients can guarantee the effect of the model in most clients. To ensure the models’ performance on exited abnormal clients and those who lack training resources, we also introduce a Federated Learning with Invariant Penalty for Generalization (FedIPG). With the assistance of the invariant penalty term, the model can achieve robust generalization capability. This approach indirectly mitigates the effects of data heterogeneity and asynchrony without additional communication overhead, making it ideal for edge intelligence systems. Our theoretical and empirical results demonstrate that FedIPG, combined with an exit strategy, enhances both in-distribution performance and out-of-distribution generalization capabilities while maintaining model convergence. This approach provides a robust framework for federated learning in resource-constrained environments while offering preliminary causal insights.
@@
Asynchronous Federated Learning, Heterogeneous environment, Out-of-Distribution Generalization, Edge intelligence
@@ I Introduction
Federated learning is a distributed computing paradigm designed to preserve privacy while enabling collaborative learning. The central tenet of FL is that data remains on local devices, with minimal interaction with a central server or other clients[1, 2]. Previous research has shown that, under ideal conditions, FL can achieve performance comparable to centralized learning. However, these ideal conditions are often challenging to attain in edge intelligence contexts, as shown in Fig.1. First, data collected from diverse sources typically exhibit significant heterogeneity and environmental interference. Second, in sophisticated scenarios involving edge intelligence, there exists a situation where communication and computational capabilities are unevenly distributed, ensuring synchronization within edge intelligence systems poses a considerable challenge. Both data heterogeneity and asynchrony can substantially degrade the performance of FL in such environments[3, 4]. While much of the existing research has focused on decrease the harm heterogeneous and asynchronous clients caused in federated aggregation, this paper introduces a novel perspective.
@@
If the performance of the trained model on aberrant clients can be guaranteed, then the withdrawal of aberrant clients from global aggregation can enhance the model’s performance on regular clients without violating the fairness principle in federated learning. Specifically, the model should be only employed on those aberrant clients after training. This exit strategy indirectly addresses the harm caused by heterogeneity and asynchrony. Our previous work has demonstrated that when latency is high, it is often more advantageous to allow computationally inefficient clients with poor communication environments to withdraw from the federated training process. In this article, we decompose the local optimal parameters of clients into two orthogonal parameter vectors: the globally effective invariant component and the locally effective environmental heterogeneous component. Through an analysis of the contribution of anomalous clients to aggregation, the effectiveness of the exit strategy is demonstrated. Although the withdrawal of anomalous clients can lead to better-performing models for In-Distribution (ID) clients, the performance of Out-Of-Distribution (OOD) clients that exit the training process is no longer guaranteed. Additionally, clients lacking essential resources for training, such as datasets or computational power, still require the model to have inference capabilities on out-of-distribution datasets. In such cases, clients may bypass training altogether and directly utilize the trained model, necessitating the model to exhibit strong OOD generalization capabilities.
The issue of OOD generalization has also been extensively studied in centralized learning [5, 6, 7]. One notable approach garnered significant attention is Invariant Risk Minimization (IRM)[8]. The core lies in decoupling data features into invariant and environmental features. Invariant features are considered stable and effective across diverse conditions, while environmental features, such as the background of an image, may introduce noise or instability[9]. For instance, in wildfire detection, flames and temperature represent invariant features. At the same time, the background (e.g., smoke or fog) serves as an environmental feature that may interfere with classification[10]. Several variants of centralized schemes have been proposed to leverage this relationship. However, there has been limited research on applying a centralized invariant scheme within the context of FL systems[11, 12].
@@
Inspired by IRM, We proposed FedIPG to realize OOD generalization in edge intelligence scenarios to guarantee the fairness of FL with the exit strategy we introduced. FedIPG employs gradients to approximate environmental heterogeneous components and constructs an invariance penalty term. This penalty term reduces the impact of environment-specific parameters, facilitating learning globally invariant parameters. This model paradigm complements the potential distribution shift and fairness violations caused by the exit strategy, ensuring that clients not participating in training have basic performance guarantees. By combining the exit strategy with FedIPG, FL can allocate more communication resources to clients with strong computational capabilities in edge learning scenarios characterized by uneven computational resource distribution and limited communication resources, thereby enabling the training of a globally invariant model. The key contributions and innovations of this paper are as follows:
@@ III-B The Profile of the Invariant Schemes and Notations
$f_{i}(w)=R_{i}(w)+\lambda|<\nabla R_{i}(w),w>|^{2}.$ (8)
Positive hyper-parameter $\lambda$ denotes the strength of invariant regularization. The quest of regularization $|<\nabla R_{i}(w),w>|^{2}$ is to find invariant features.
Significantly, each client’s update still optimizes towards its local optimal parameter. Due to the invariance penalty term altering the distribution of the client loss functions, the clients exhibit a smoother landscape at the invariance optimal parameter point, where $\|<\nabla R(w),w>\|^{2}=0$. After weighted summation at the server, the global optimal parameter $w^{*}$ can be considered the midpoint of the client optimal parameters $w^{*}_{i}$. Guided by the penalty term, the summation yields a new invariance optimal parameter $w^{I}$. $w^{I}$ exhibits relatively smooth gradients across any client, indicating a focus on invariance features while potentially neglecting rapidly changing environmental information.
@@ IV-A Theoretical Measure for Clients’ Contribution
In this section, we define the contribution of a client’s participation in a round of aggregation. When there is a probability that a client’s contribution is negative, it implies that the client’s participation may have a detrimental effect on the aggregation. By orthogonally decomposing parameters, the contributions of heterogeneous and asynchronous clients are quantified. The probability of an aberrant client’s contribution being negative grows with the size of the heterogeneous component and the delay, suggesting that excluding such clients benefits aggregation.
@@ IV-B Contribution Analysis for Aberrant clients
Here, we assume the presence of a highly heterogeneous client $i=h$, while the data of other clients can be approximated as independent and identically distributed (IID), i.e., $U_{i}\approx V_{i}\approx 1$. Under this assumption, the upper bound of $C_{i}^{t}$ can be expressed as:
@@
When this external heterogeneous client $h$ participates in training, whose optimal parameter is $w^{*}_{h}=w_{I}+w^{e}_{h}$, we obtain the following conclusion:
@@
For a optimization started from the origin of coordinates, while $w^{t}=\kappa\frac{w^{I}}{\|w^{I}\|}+\nu\frac{w^{e}_{h}}{\|w^{e}_{h}\|}$, $\kappa\in(0,1)$ and $\nu\in(-\epsilon,\epsilon)$, the contribution of client $h$ in iteration $t$ can be represent as:
$C^{t}_{h}=\eta\frac{n_{i}}{n}(\frac{U_{i}((1-\kappa)^{2}-\nu(\|w^{e}_{h}\|-\nu))}{\sqrt{(1-\kappa)^{2}+\nu^{2}}\sqrt{(1-\kappa)^{2}+(\|w^{e}_{h}\|-\nu)^{2}}}-\frac{L}{2}V_{i}^{2})$ (15)
@@
Under the condition of sufficient data volume, the impact of a single client on global aggregation can be considered minimal, which means $\epsilon<\|w^{e}_{h}\|$. When the magnitude of $\|w^{I}\|$ is fixed, the denominator of (15) increases as $\|w^{e}_{h}\|$ grows. For $\nu\leq 0$, the numerator decreases with the increase of $\|w^{e}_{h}\|$. For $\nu\geq 0$, the numerator increases with $\|w^{e}_{h}\|$, but at a slower rate than the denominator. Clearly, as data heterogeneity increases, the value of client participation in global aggregation diminishes. Clients with strong heterogeneity are more likely to have a negative impact.
Next, we analyze the impact of delay $\tau$ to $P^{t}_{i}$. To simplify the problem, we assume all clients include highly homogeneous data, which means $\nabla R_{i}(w)$ is equal to $\nabla R(w)$. The upper bound of $C_{i}^{t}$ can be expressed as:
@@
$=\eta\frac{n_{i}}{n}\frac{\|\nabla R(w^{t-\tau_{i}^{t}})\|}{\|\nabla R(w^{t})\|^{2}}-\frac{L}{2}\frac{\|\nabla R_{i}(w^{t-\tau_{i}^{t}})\|^{2}}{\|\nabla R(w^{t})\|^{2}}$ (16)
Notably, $C_{i}^{t}$ degenerates into a function of $\frac{\|\nabla R(w^{t-\tau_{i}^{t}})\|}{\|\nabla R(w^{t})\|}$. When $\tau_{i}^{t}=0$, $C_{i}^{t}>0$, which implies $L<2$, a condition that typically holds under normal circumstances.
To analyze the upper bound, the squared norm of the gradients is uniformly bounded[38].
@@
Admittedly, whether for heterogeneous clients or asynchronous clients, we are discussing the worst-case scenario here, and the lower bound of $C_{i}^{t}$ may not necessarily be reached. However, when data resources are sufficient, prohibiting such clients from participating can effectively mitigate this risk.
@@ V-A Out-of-Distributed Generalization by Invariant Penalty Term
In (20), $\mathcal{R}_{i}(0)-\mathcal{R}_{j}(0)$ is an inherent property of client data that does not change with change of parameters. So, according to Equation (20), by minimizing $<\nabla R(w),w>$ in the last term, the parameter with better performance on the generalized dataset can be found. Additionally, as seen in Equation (20), the maximum upper bound is related to $R_{i}(0)-R_{j}(0)$, the maximum difference in the client loss functions at the zero point. The exit strategy for heterogeneous clients reduces $R_{i}(0)-R_{j}(0)$, thereby significantly lowering this upper bound.
@@ V-B Convergence Analysis of Invariant Asynchronous Federated Learning Algorithm
This implies that when designing the loss function, careful consideration must be given to the value of $\lambda$. Indeed, due to the inherent difficulty in assessing the convexity and smoothness of the empirical risk loss, it is practically impossible to quantitatively compute an upper bound for $\lambda$. However, in practical applications, we observe that a relatively small $\lambda$ typically suffices to ensure the convergence of the algorithm, indicating that ensuring this upper bound is not particularly challenging. Next, we will demonstrate that under this design of $\lambda$, the loss function incorporating the penalty term will simultaneously exhibit smoothness.
@@
Based on the above Lemma, the convergence of loss function FedIPM without isolation of data can be proved in Appendix-B Based on the above conclusions, we have established that the loss function incorporating the penalty term is convex and smooth. Additionally, in the centralized setting, grouping data by environment allows us to obtain the optimal invariant parameters under the current penalty term using gradient descent. Next, we will shift our focus to the context of edge intelligence and further analyze the convergence properties of invariant learning within a federated framework.
In the server aggregation convergence analysis, we assume all selected $\mathcal{I}_{ID}$ devices are able to participate in every aggregation steps. For subsequent proof of convergence, heterogeneity measurement between environments is defined as:
@@ VII Conclusion
Appendix A: Proof of Theorem 1 and 2
Theorem 1.
(The Value Analysis for Heterogeneous client)
For a optimization started from the origin of coordinates, while $w^{t}=\kappa\frac{w^{I}}{\|w^{I}\|}+\nu\frac{w^{e}_{h}}{\|w^{e}_{h}\|}$, $\kappa\in(0,1)$ and $\nu\in(-\|w^{e}_{h}\|,\|w^{e}_{h}\|)$. The value of client $h$ in iteration $t$ can be represent as:
$p^{t}_{h}=\eta\frac{n_{i}}{n}(\frac{U_{i}((1-\kappa)^{2}-\nu(\|w^{e}_{h}\|-\nu))}{\sqrt{(1-\kappa)^{2}+\nu^{2}}\sqrt{(1-\kappa)^{2}+(\|w^{e}_{h}\|-\nu)^{2}}}-\frac{L}{2}V_{i}^{2})$ (55)
Proof:
$\theta_{i}^{t}$ represents the angle between $\nabla R_{i}(w^{t})$ and $\nabla R(w^{t})$, we use $w^{t}-w^{*}$ and $w^{t}-w^{*}_{e}$ to approximate $\nabla R(w^{t})$ $\nabla R_{i}(w^{t})$, which means $\cos{\theta_{i}^{t}}$ can be written as
$\cos{\theta_{i}^{t}}=<\nabla R(w^{t}),\nabla R_{i}(w^{t})>$
$\approx\frac{<w^{t}-w^{I},w^{t}-w^{*}_{h}>}{\|w^{t}-w^{I}\|\|w^{t}-w^{*}_{h}\|}$
$=\frac{<\kappa\frac{w^{I}}{\|w^{I}\|}+\nu\frac{w^{e}_{h}}{\|w^{e}_{h}\|}-w^{I},\kappa\frac{w^{I}}{\|w^{I}\|}+\nu\frac{w^{e}_{h}}{\|w^{e}_{h}\|}-w^{*}_{h}>}{\|\frac{w^{I}}{\|w^{I}\|}+\nu\frac{w^{e}_{h}}{\|w^{e}_{h}\|}-w^{I}\|\|\frac{w^{I}}{\|w^{I}\|}+\nu\frac{w^{e}_{h}}{\|w^{e}_{h}\|}-w^{*}_{h}\|}$
$=\frac{(\|w^{I}\|-\kappa)^{2}-\nu(\|w^{e}_{h}\|-\nu)}{(\|w^{I}\|-\kappa)^{2}+(\|w^{e}_{h}\|-\nu)^{2}}$ (56)
we introduce (56) into (14), we have
$p^{t}_{h}=\eta\frac{n_{i}}{n}(\frac{U_{i}((1-\kappa)^{2}-\nu(\|w^{e}_{h}\|-\nu))}{\sqrt{(1-\kappa)^{2}+\nu^{2}}\sqrt{(1-\kappa)^{2}+(\|w^{e}_{h}\|-\nu)^{2}}}-\frac{L}{2}V_{i}^{2})$ (57)
Appendix B: Centralized convergence of FedIPG
Based on the above Lemma, the convergence of loss function FedIPM without isolation of data will be proved as the benchmark of federated framework. The weighted summation of a convex function is a convexity preserving operation. Based on (28), we obtain:
Theorem 2.
If gradient descent for $t$ iterations is performed with a fixed step size $\eta\leq\frac{1}{L^{\prime}}$, it will yield a solution $f(w^{T})$ which satisfies
$f(w^{T})-f(w^{I})\leq\frac{\|w^{0}-w^{I}\|^{2}_{2}}{2T\eta},$ (58)
where $w^{I}$ is the optimal invariant parameter and $\eta$ is the learning rate. If the optimization started from the origin of the parameter space,, then:
$f(w^{T})-f(w^{I})\leq\frac{\|w^{I}\|^{2}_{2}}{2T\eta}.$ (59)
Proof.
It can deduced that global loss function $f(w)$ also has the property of smoothness. If we assume $\eta\leq\frac{1}{L^{\prime}}$, we can conclude the following:
$f(w^{t+1})\leq f(w^{t})+\frac{1}{2}\eta\|\nabla f(w^{t})\|^{2}_{2}.$ (60)
Since weighted summation is a operation that preserve convexity, $f(w)$ is a convex function whose optimal parameter is $w=w^{I}$. We can write
$f(w^{I})\geq f(w^{t})+\nabla f(w^{t})^{\top}(w^{I}-w^{t}).$ (61)
Plugging above inequality to the in, we obtain:
$f(w^{t+1})-f(w^{I})$ (62)
$\leq$ $\frac{1}{2\eta}[2\eta\nabla f(w^{t})^{\top}(w^{t}-w^{I})-\eta^{2}\|\nabla f(w^{t})\|^{2}_{2})]$
$\leq$ $\frac{1}{2\eta}(\|w^{t}-w^{I}\|^{2}_{2}-\|w^{t+1}-w^{I}\|).$
Summing over all iterations, since $f(w)$ is decreasing on every iteration, if the optimization started from the origin of coordinates, we conclude that:
$f(w^{t})-f(w^{I})$ $\leq\frac{1}{t}\sum^{t}_{i=1}(f(w^{i})-f(w^{I})$ (63)
$\leq\frac{\|w^{0}-w^{I}\|^{2}_{2}}{2\eta t}$
$=\frac{\|w^{I}\|^{2}_{2}}{2\eta t}.$
Gradient decent is guaranteed to converge at $w^{I}$ when loss function is calculated without isolation of data. ∎
Appendix C: Proof of Theorem 6
Theorem 3.
(The Upper Bound of Loss Functions with Invariant Penalty)
With the invariant penalty, the loss function $f_{(}w)$ holds
Proof:
Consider Theorem 4, we have
$f(w^{t})-f(w^{I})\leq<\nabla f(w^{t}),w^{t}-w^{I}>-\frac{\mu^{\prime}}{2}\|w^{I}-w^{t}\|^{2}$
$\leq<\nabla f(w^{t}),w^{t+1}-w^{*}>+<\nabla f(w^{t}),w^{t}-w^{t+1}>$
Based on Theorem 5, the upper bound of $<\nabla f(w^{t}),w^{t}-w^{t+1}>$ can be introduced into the above inequality, which turn to
$f(w^{t})-f(w^{I})\leq<\nabla f(w^{t}),w^{t+1}-w^{*}>+f(w^{t})$ (66)
$-f(w^{t+1})+\frac{L^{\prime}}{2}\|w^{t+1}-w^{t}\|^{2}$ (67)
By applying rearrangement to above inequality, we obtain
Since we have specified the update policy in Equation (2), we have $\sum_{i\in\mathcal{I}_{ID}}a_{i}\nabla f_{i}(w^{t})=-\frac{1}{\eta}(w^{t+1}-w^{t})$. The above inequality can be rewrite as
$f(w^{t+1})-f(w^{I})$
$\leq<\sum_{i\in\mathcal{I}_{ID}}a_{i}\nabla f_{i}(w^{t}),w^{t+1}-w^{*}>+\frac{L^{\prime}}{2}\|w^{t+1}-w^{t}\|^{2}$
$\leq-\frac{1}{\eta}<w^{t+1}-w^{t},w^{t+1}-w^{*}>+\frac{L^{\prime}}{2}\|w^{t+1}-w^{t}\|^{2}$ (69)
$f(w^{t+1})-f(w^{I})$
$\leq-\frac{1}{2\eta}(\|w^{t+1}-w^{t}\|^{2}+\|w^{t+1}-w^{I}\|^{2})+\frac{L^{\prime}}{2}\|w^{t+1}-w^{t}\|^{2}$
$=\frac{1}{2\eta}(\|w^{t+1}-w^{t}\|^{2}+2\|w^{t+1}-w^{I}\|^{2})+(L^{\prime}-\frac{1}{\eta})\|w^{t+1}-w^{t}\|^{2}$ (70)
According to the inequality of arithmetic and geometric means, we have $\|w^{t+1}-w^{I}\|^{2}\leq 2\|w^{t}-w^{I}\|^{2}+2\|w^{t+1}-w^{I}\|^{2}$. Furthermore, we have
$+\frac{1}{2}(L^{\prime}-\frac{1}{\eta})\|w^{t+1}-w^{t}\|^{2}$ (71)
By summing (71) over every training iterations, we have
$\sum^{T}_{t=1}[f(w^{t+1})-f(w^{I})]$ $\leq\frac{1}{2\eta}(\|w^{1}-w^{I}\|^{2}-\|w^{T+1}-w^{I}\|^{2})$
$+\frac{1}{2}(L^{\prime}-\frac{1}{\eta})\sum^{T}_{t=1}\|w^{t+1}-w^{t}\|^{2}$ (72)
Due to the inequality of arithmetic and geometric means, we have
$=\frac{1}{T}\|w^{T+1}-w^{1}\|^{2}$ (73)
Based on above inequalities, we get
$\sum^{T}_{t=1}[f(w^{t+1})-f(w^{I})]$
$\leq\frac{1}{2\eta}(\|w^{1}-w^{I}\|^{2}-\|w^{T+1}-w^{I}\|^{2})$
$-\frac{1}{2T}(\frac{1}{\eta}-L^{\prime})\|w^{T+1}-w^{1}\|^{2}$ (74)
$\leq\frac{B^{2}}{2\eta}+\frac{L^{\prime}}{2T}\|w^{T+1}-w^{1}\|^{2}$ (75)
According to (28). we have
$f(\hat{w}(T))-f(w^{I})\leq\frac{1}{T}\sum^{T}_{t=1}[f(w^{t+1})-f(w^{I})]$ (76)
Since $\nabla f_{i}(w^{*}_{i})=0$ has been proved in (36), we have
$\|w^{t}-w^{*}_{i}\|^{2}\leq\frac{1}{\mu}[f_{i}(w^{t})-f_{i}(w^{*}_{i})]$ (78)
So $\|w^{t}-w^{I}\|^{2}$ can be limited by following elements
$\|w^{t}-w^{I}\|^{2}$ $=\sum_{i\in\mathcal{I}_{ID}}a_{i}\|w^{t}-w\|^{2}$
Hence, $\|w^{T+1}-w^{1}\|^{2}$ can be bounded as
$\|w^{T+1}-w^{1}\|^{2}\leq$ $\frac{4}{\mu^{\prime}}f(w^{T+1})+\frac{4}{\mu}f(w^{1})$
$+\sum_{i\in\mathcal{I}_{ID}}a_{i}[4\|w^{I}-w_{i}^{*}\|^{2}-\frac{8}{\mu^{\prime}}f_{i}(w^{*}_{i})]f_{i}(w^{*}_{i}))$
$\leq\frac{4L^{\prime}B^{2}}{\mu^{\prime}}+\sum_{i\in\mathcal{I}_{ID}}4a_{i}\|w^{I}-w_{i}^{*}\|^{2}$
$+\sum_{i\in\mathcal{I}_{ID}}\frac{8a_{i}}{\mu^{\prime}}(f_{i}(w^{I})-f_{i}(w^{*}_{i}))$
$\leq\frac{4L^{\prime}B^{2}}{\mu^{\prime}}+\sum_{i\in\mathcal{I}_{ID}}a_{i}(4+\frac{4L^{\prime}}{\mu^{\prime}})\|w^{I}-w_{i}^{*}\|^{2}$
$\leq\frac{4L^{\prime}B^{2}}{\mu^{\prime}}+(4+\frac{4L^{\prime}}{\mu^{\prime}})\phi^{2}$ (80)
Plugging (2) in (75), we have
According to Theorem 2 (28), we have
$f(\hat{w}(T))-f(w^{*})$ $\leq\frac{1}{T}\sum^{T}_{t=1}[f(w^{t+1})-f(w^{I})]$

### v2

# Invariant Federated Learning for Edge Intelligence: Mitigating Heterogeneity and Asynchrony via Exit Strategy and Invariant Penalty

@@ Abstract
This paper provides an invariant federated learning system for resource-constrained edge intelligence. This framework can avoid the impact of heterogeneity and asynchrony by exit strategy and invariant penalty. We decompose local information into two orthogonal components to measure the contribution or impact of heterogeneous and asynchronous clients. We propose that the exit of abnormal clients can guarantee the effect of the model on most clients. Meanwhile, to ensure the models’ performance on exited abnormal clients and those who lack training resources, we propose Federated Learning with Invariant Penalty for Generalization (FedIPG) based on the invariant orthogonal decomposition of parameters. Theoretical proof shows that FedIPG reduces the Out-Of-Distribution prediction loss without increasing the communication burden. The performance of FedIPG combined with an exit strategy is tested empirically in multiple scales using four datasets. It shows our system can enhance In-Distribution performance and outperform the state-of-the-art algorithm in Out-Of-Distribution generalization while maintaining model convergence. Additionally, the results of the visual experiment prove that FedIPG contains preliminary causality in terms of ignoring confounding features.
@@
Invariant learning, Asynchronous Federated Learning, Heterogeneous environment, Out-of-Distribution Generalization, Edge intelligence
@@ I Introduction
Federated learning is a distributed computing paradigm designed to preserve privacy while enabling collaborative learning. The central tenet of FL is that data remains on local devices, with minimal interaction with a central server or other clients[1, 2]. Previous research has shown that, under ideal conditions, FL can achieve performance comparable to centralized learning. However, these ideal conditions are often challenging to attain in edge intelligence contexts, as shown in Fig.1. First, data collected from diverse sources typically exhibit significant heterogeneity and environmental interference. Second, in sophisticated scenarios involving edge intelligence, there exists a situation where communication and computational capabilities are unevenly distributed, ensuring synchronization within edge intelligence systems poses a considerable challenge. Both data heterogeneity and asynchrony can substantially degrade the performance of FL in such environments[3, 4].
@@
While much of the existing research has focused on decreasing the harm heterogeneous and asynchronous clients caused in federated aggregation, inspired by invariant learning, we introduce a novel perspective that suggests heterogeneous and asynchronous clients exit the training process. Specifically, the model should be only employed on those aberrant clients after training. Our previous work has demonstrated that when latency is high, it is often more advantageous to allow computationally inefficient clients with poor communication environments to withdraw from the federated training process[5]. In this article, we decompose the local optimal parameters of clients into two orthogonal parameter vectors: the globally practical invariant component and the locally effective environmental heterogeneous component[6]. We proved that, as the size of the heterogeneous component grows and the delay increases, the anomalous clients are more likely to provide a negative effect rather than a positive contribution, which means the exit strategy can protect the model from disturbance by heterogeneous and asynchronous clients. Suppose models’ generalization capability can guarantee the performance on aberrant clients’ prediction. In that case, the withdrawal of aberrant clients from global aggregation can enhance the model’s performance on regular clients, which account for the majority, without harming the rights of the few exited clients. Although the withdrawal of anomalous clients indirectly addresses the harm caused by heterogeneity and asynchrony, leading to better-performing models for In-Distribution (ID) clients, the performance of Out-Of-Distribution (OOD) clients that exit the training process is no longer guaranteed. Additionally, clients lacking essential resources for training, such as datasets or computational power, still require the model to have inference capabilities on out-of-distribution datasets. In such cases, clients may bypass training altogether and directly utilize the trained model, necessitating the model to exhibit strong OOD generalization capabilities. The issue of OOD generalization has also been extensively studied in centralized learning [7, 8, 9]. One notable approach garnered significant attention is Invariant Learning[10]. The core lies in decoupling data features into invariant and environmental features. Invariant features are considered stable and effective across diverse conditions, while environmental features, such as the background of an image, may introduce noise or instability[11]. For instance, in wildfire detection, flames and temperature represent invariant features. At the same time, the background (e.g., smoke or fog) serves as an environmental feature that may interfere with classification[12]. Several variants of centralized schemes have been proposed to leverage this relationship. However, there has been limited research on applying a centralized invariant scheme within the context of FL systems[13, 14].
@@
Inspired by Invariant Risk Minimization[15], We proposed FedIPG to realize OOD generalization in edge intelligence scenarios to satisfy the prediction needs of exited clients in the exit strategy we proposed. FedIPG employs gradients to approximate environmental heterogeneous components and constructs an invariance penalty term. This penalty term reduces the impact of environment-specific parameters, facilitating learning globally invariant parameters. This model paradigm complements the potential distribution shift and fairness violations caused by the exit strategy, ensuring that clients not participating in training have basic performance guarantees. By combining the exit strategy with FedIPG, FL can allocate more communication resources to clients with strong computational capabilities in edge learning scenarios characterized by uneven computational resource distribution and limited communication resources, thereby enabling the training of a globally invariant model. The key contributions and innovations of this paper are as follows:
@@ III-B The Profile of the Invariant Schemes and Notations
$f_{i}(w)=R_{i}(w)+\lambda\|<\nabla R_{i}(w),w>\|^{2}.$ (8)
Positive hyper-parameter $\lambda$ denotes the strength of invariant regularization. The quest of regularization $\|<\nabla R_{i}(w),w>\|^{2}$ is to find invariant features.
Significantly, each client’s update still optimizes towards its local optimal parameter. Due to the invariance penalty term altering the distribution of the client loss functions, the clients exhibit a smoother landscape at the invariance optimal parameter point, where $<\nabla R(w),w>=0$. After weighted summation at the server, the global optimal parameter $w^{*}$ can be considered the midpoint of the client optimal parameters $w^{*}_{i}$. Guided by the penalty term, the summation yields a new invariance optimal parameter $w^{I}$. $w^{I}$ exhibits relatively smooth gradients across any client, indicating a focus on invariance features while potentially neglecting rapidly changing environmental information.
@@ IV-A Theoretical Measure for Clients’ Contribution
In this section, we define the contribution of a client’s participation in a round of aggregation. When there is a probability that a client’s contribution is negative, it implies that the client’s participation may have a detrimental effect on the model’s performance. By orthogonally decomposing parameters, the contributions of heterogeneous and asynchronous clients are quantified.
@@ IV-B Contribution Analysis for Aberrant clients
Here, we assume the presence of a highly heterogeneous client $i$, while the data of other clients can be approximated as independent and identically distributed (IID), i.e., $U_{i}\approx V_{i}\approx 1$. Under this assumption, the upper bound of $C_{i}^{t}$ can be expressed as:
@@
When this external heterogeneous client $i$ participates in training, whose optimal parameter is $w^{*}_{i}=w_{I}+w^{e}_{i}$, we obtain the following conclusion:
@@
For a optimization started from the origin of coordinates, while $w^{t}=\kappa\frac{w^{I}}{\|w^{I}\|}+\nu\frac{w^{e}_{i}}{\|w^{e}_{i}\|}$, $\kappa\in(0,1)$ and $\nu\in(-\epsilon,\epsilon)$, the contribution of client $i$ in iteration $t$ can be represent as:
$C^{t}_{i}=\eta\frac{n_{i}}{n}(\frac{U_{i}((1-\kappa)^{2}-\nu(\|w^{e}_{i}\|-\nu))}{\sqrt{(1-\kappa)^{2}+\nu^{2}}\sqrt{(1-\kappa)^{2}+(\|w^{e}_{i}\|-\nu)^{2}}}-\frac{L}{2}V_{i}^{2})$ (15)
@@
Under the condition of sufficient data volume, the impact of a single client on global aggregation can be considered minimal, which means $\epsilon<\|w^{e}_{i}\|$. When the magnitude of $\|w^{I}\|$ is fixed, the denominator of (15) increases as $\|w^{e}_{i}\|$ grows. For $\nu\leq 0$, the numerator decreases with the increase of $\|w^{e}_{i}\|$. For $\nu\geq 0$, the numerator increases with $\|w^{e}_{i}\|$, but at a slower rate than the denominator. Clearly, as data heterogeneity increases, the value of client participation in global aggregation diminishes. Clients with strong heterogeneity are more likely to have a negative impact.
Next, we analyze the impact of delay $\tau$ to $C^{t}_{i}$. To simplify the problem, we assume all clients include highly homogeneous data, which means $\nabla R_{i}(w)$ is equal to $\nabla R(w)$. The upper bound of $C_{i}^{t}$ can be expressed as:
@@
$=\eta\frac{n_{i}}{n}\frac{\|\nabla R(w^{t-\tau_{i}^{t}})\|}{\|\nabla R(w^{t})\|}-\frac{L}{2}\frac{\|\nabla R_{i}(w^{t-\tau_{i}^{t}})\|^{2}}{\|\nabla R(w^{t})\|^{2}}$ (16)
When $\tau_{i}^{t}=0$, $C_{i}^{t}>0$, which implies $L<2$, a condition that typically holds under normal circumstances. Notably, $C_{i}^{t}$ degenerates into a function of $\frac{\|\nabla R(w^{t-\tau_{i}^{t}})\|}{\|\nabla R(w^{t})\|}$. We next analyze the trend of $\frac{\|\nabla R(w^{t-\tau_{i}^{t}})\|}{\|\nabla R(w^{t})\|}$ as $\tau$ changes, which indirectly reflects the variation in client $C^{t}_{i}$. Most importantly, we examine the negativity of $C^{t}_{i}$ when $\tau$ is large. The squared norm of the gradients is uniformly bounded[41].
@@
As a result, the probability of an aberrant client’s contribution being negative grows with the size of the heterogeneous component and the delay, suggesting that excluding such clients benefits aggregation. Prohibiting aberrant clients from participating can effectively mitigate this negetive contribution.
@@ V-A Out-of-Distributed Generalization by Invariant Penalty Term
In Equation (20), $\mathcal{R}_{i}(0)-\mathcal{R}_{j}(0)$ is an inherent property of client data that does not change with change of parameters. So, by minimizing $<\nabla R(w),w>$ in the last term, the parameter with better performance on the generalized dataset can be found. Additionally, as seen in Equation (20), the maximum upper bound is related to $R_{i}(0)-R_{j}(0)$, the maximum difference in the client loss functions at the zero point. The exit strategy for heterogeneous clients reduces $R_{i}(0)-R_{j}(0)$, thereby significantly lowering this upper bound.
@@ V-B Convergence Analysis of Invariant Asynchronous Federated Learning Algorithm
This implies that when designing the loss function, careful consideration must be given to the value of $\lambda$. Indeed, due to the inherent difficulty in assessing the convexity and smoothness of the empirical risk loss, it is practically impossible to quantitatively compute an upper bound for hyperparameter $\lambda$. However, in numerical simulation, we observe that a relatively small $\lambda$ typically suffices to ensure the convergence of the algorithm, indicating that ensuring this upper bound is not particularly challenging. Next, we will demonstrate that under this design of $\lambda$, the loss function incorporating the penalty term will simultaneously exhibit smoothness.
@@
Based on the above Lemma, the convergence of loss function FedIPM without isolation of data can be proved in Appendix-B Based on the above conclusions, we have established that the loss function incorporating the penalty term is convex and smooth. Additionally, in the centralized setting, grouping data by environment allows us to obtain the optimal invariant parameters under the current penalty term using gradient descent. Next, we will shift our focus to the context of edge intelligence, analyzing the convergence properties of invariant learning within a federated framework.
In the server aggregation convergence analysis, we assume all selected $\mathcal{I}_{ID}$ devices are able to participate in every aggregation steps. Based on the exit strategy discussed earlier, we posit that there exists an upper bound on the heterogeneity differences among all participating clients in the training process. This is consistent with the assumptions commonly made in FL researches[46].